How do you solve -3(s+2)= -2(s+4)-s?
Thank you

Answer by unlockmath(1688)

Hello,
This -3(s+2)= -2(s+4)-s can be rewritten as:
-3s-6= -2s-8-s Combine like terms will be and add 6 to both sides:
-3s=-3s-2 Add 3s to both sides:
0=-2 As you can see there is no solution to this problem since we know 0 does not equal -2
Make sense?
RJ Check out a new book at:
www.math-unlock.com